# Gang Li244footnotemark: 4     Mingcheng Li144footnotemark: 4     Shuaibing Wang144footnotemark: 4     Jiawei Chen144footnotemark: 4     Yue Jiang144footnotemark: 4 Qingyao Xu144footnotemark: 4     Ke Li244footnotemark: 4     Peng Zhai1,311footnotemark: 1     Lihua Zhang1,3,4,511footnotemark: 1

### v2

@@ Abstract
Developing intelligent pediatric consultation systems offers promising prospects for improving diagnostic efficiency, especially in China, where healthcare resources are scarce. Despite recent advances in Large Language Models (LLMs) for Chinese medicine, their performance is sub-optimal in pediatric applications due to inadequate instruction data and vulnerable training procedures. To address the above issues, this paper builds PedCorpus, a high-quality dataset of over 300,000 multi-task instructions from pediatric textbooks, guidelines, and knowledge graph resources to fulfil diverse diagnostic demands. Upon well-designed PedCorpus, we propose PediatricsGPT, the first Chinese pediatric LLM assistant built on a systematic and robust training pipeline. In the continuous pre-training phase, we introduce a hybrid instruction pre-training mechanism to mitigate the internal-injected knowledge inconsistency of LLMs for medical domain adaptation. Immediately, the full-parameter Supervised Fine-Tuning (SFT) is utilized to incorporate the general medical knowledge schema into the models. After that, we devise a direct following preference optimization to enhance the generation of pediatrician-like humanistic responses. In the parameter-efficient secondary SFT phase, a mixture of universal-specific experts strategy is presented to resolve the competency conflict between medical generalist and pediatric expertise mastery. Extensive results based on the metrics, GPT-4, and doctor evaluations on distinct doctor downstream tasks show that PediatricsGPT consistently outperforms previous Chinese medical LLMs. Our model and dataset will be open-source for community development.
@@ 2 Related Work
Chinese Large Language Model Evolution. The emergence of Large Language Models (LLMs) dominated by ChatGPT [31] and GPT-4 [5] has revolutionized the paradigm for novel human-machine interaction. Driven by learning-oriented technologies [11, 12, 13, 50, 51, 52], pragmatic instruction [30, 45] and preference optimization [7, 32] strategies enable LLMs to address complex generation tasks with aligned human intentions. Despite improvements, large-scale resources for training general LLMs [26, 41, 42] are anchored in the English corpora, limiting their abilities to respond reliably in extensive Chinese application scenarios. Recently, researchers [17, 56] have attempted to enhance the comprehension and execution of Chinese instructions in open-source LLMs by augmenting Chinese vocabulary and data (e.g., Chinese LLaMA and Alpaca [17]). To facilitate Chinese-specific demands, several LLMs trained from scratch exhibit remarkable Chinese proficiency due to multilingual data resources, such as the Baichuan [2, 49], General Language Model (GLM) [19, 54], and Qwen [6] families. In this work, the Baichuan2-Base series is utilized as the foundation model for our PediatricsGPT, given its comprehensive potential among similar contenders.
@@ 3.1 PedCorpus: Multi-task Medical Instruction Dataset
Specialized Pediatric Data. Extracting pediatric data from textbooks, guidelines, and knowledge graphs ensures knowledge professionalism. Specifically, we automatically extract standard medical definitions and descriptions from physical textbooks covering 131 disease types in 11 broad categories. Over 500 corresponding disease guidelines are collected, including diagnostic protocols and treatment consensus. Additionally, extensive knowledge entities are sampled from ternary instances in the knowledge graphs. Based on complementary resources, we introduce a role-playing-driven instruction building rule via GPT-4 API that produces well-organized instructions to enable accurate and humanistic model responses. The detailed building procedure is shown in Appendix A.1.
@@ 3.2 Hybrid Instruction Pre-training in CPT
Continuous Pre-Training (CPT) is essential in developing domain-specific models [14, 46, 53] since it can break the scaling law [22] to a certain extent. For this purpose, we introduce the PedCorpus-CPT dataset to ensure a high-quality pre-training corpus. From Table 1, PedCorpus-CPT consists of three-part data components. (i) We integrate plain texts from vanilla pediatric textbooks, guidelines, and knowledge graphs. (ii) The filtered Chinese Wikipedia [3] is also considered to achieve the model’s trade-off for medical-general knowledge memory capacity. (iii) In practice, we observe that CPT leads to catastrophic forgetting of the models at follow-up due to different data distribution and format discrepancies compared to the original pre-training and SFT. Thus, we introduce a hybrid instruction pre-training mechanism to bridge these discrepancies. The core philosophy is to assemble instruction data from PedCorpus with Input-Output forms into Completion forms, which are then assimilated into plain texts to provide multi-task and complementary information. Our mechanism effectively mitigates inconsistencies between the internal-injected medical knowledge of the foundation model while reinforcing medical domain adaptation. Moreover, we take PedCorpus as the seed instructions to improve multiple-department corpus density and breadth via knowledge-enhanced prompts. The prompt template is shown in Appendix B.
@@ 4.2 Model Zoo
We compare a series of LLMs for comprehensive evaluations. Concretely, Baichuan2-7B/13B (Chat) models [49] are trained on 2.6 trillion tokens as the baselines, which have excellent abilities in different domains. For reproducible Chinese medical works, DISC-MedLLM (13B) [8] is fine-tuned through reconstructed medical dialogues and behavioural preference instructions. HuatuoGPT (13B) [55] performs SFT based on mixed instruction data and introduce human feedback in RLHF. HuatuoGPT-II (13B) [14] enhances the medical-specific domain adaptation of LLMs through one-stage unified training. Zhongjing (13B) [53] implements a complete pipeline based on Ziya-LLaMA-13B to enhance the model’s multi-turn medical conversation abilities. ChatGPT [31] has impressive performance in general medical fields as a closed-source model developed by OpenAI.
@@ 4.4 Ablation Studies
Necessity of Supervised Fine-tuning. (i) We observe consistent performance gaps when removing the Full-parameter SFT (FSFT) and Parameter-efficient SFT (PSFT) phases, respectively. This makes sense since SFTs are necessary to activate the model’s healthcare instruction-following capabilities. (ii) Moreover, PSFT is more critical for three pediatric applications because it facilitates pediatric-related knowledge accumulation, while FSFT focuses on consolidating general medical representations. (iii) Then, we replace the proposed Mixture of Universal-specific Experts (MUE) version with the vanilla single LoRA. The reduced performance on pediatric EviDiag and TreRecom benchmarks verifies that it is essential to introduce multiple LoRAs that act as specific experts on different tasks. A reasonable explanation is that the single-LoRA model suffers from the task competition between learning the knowledge question-answer and mastering the diagnostic recommendation abilities. (iv) Furthermore, we find that the universal LoRA expert significantly improves the results on the general medical benchmarks (i.e., CMD and webMedQA), proving that it mitigates the competency conflict between general medical and pediatric knowledge.
@@ 4.6 Visualization Analysis of Model Responses
To intuitively compare the output quality of medical LLMs, we show the responses of different models for each of the three types of medical inquiries in Figure 7 and Figures 14&15 from Appendix E. From the results, Zhongjing offers insufficient information due to limited output content. Although HuatuoGPT-II gives well-organized responses, it lacks accuracy and informativeness. In comparison, our model can provide more specialized and detailed medical knowledge and diagnostic guidance in extended response contexts, confirming its application potential in diverse healthcare services.
5 Conclusion
This paper presents PediatricsGPT, a Chinese medical LLM assistant with medical generalist and pediatric expertise capabilities. Based on the well-designed PedCorpus dataset, PediatricsGPT undergoes a systematic and robust procedure ranging from continuous pre-training and supervised fine-tuning to human preference optimization, leading to competence in different pediatric and general healthcare service scenarios. Extensive experimental results under multi-dimensional evaluation patterns demonstrate that our model outperforms currently available Chinese medical LLMs, providing a potential solution for promoting reliable and intelligent interactive diagnosis and treatment. We provide more discussions of broader impacts, limitations, and future work in Appendices F, G, and H.
Acknowledgment. This work is supported in part by the National Key R&D Program of China under Grant 2021ZD0113503, in part by the Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0103, and in part by the Shanghai Municipal Science and Technology Committee of Shanghai Outstanding Academic Leaders Plan (No. 21XD1430300).
@@ A.1 Role-playing-driven Instruction Building Rule
After integrating pediatric textbooks, guidelines, and knowledge graphs into a consolidated textual database, the content is segmented according to individual diseases. Subsequently, two instances of the GPT-4 model are deployed, designated as the “inquirer” and the “expert pediatrician” respectively. Disease-specific segments are then fed into the “inquirer” GPT-4, tasked with formulating a series of relevant and scholarly pediatric inquiries. Following this, the original disease segments and the formulated inquiries are fed into the “expert pediatrician” GPT-4 to generate precise responses for each inquiry, leveraging the segmented text as the contextual reference. We show the prompt templates for the “inquirer” and “expert pediatrician” in Figures 8 and 9, respectively.
@@ A.2 Vanilla Doctor-patient Conversation Regularization
We guide GPT-4 to regularize concise and noisy doctor responses in authentic doctor-patient consultations by the context learning strategy. Specifically, we manually craft 100 instruction examples with high-quality content to allow GPT-4 to learn doctor-like and patient-friendly behavioural styles. In each round of regularization, we randomly sample 10 out of 100 examples to perform 10-shot context prompts. Immediately after that, vanilla dialogues are fed to GPT-4 as seed instructions to optimize instructions according to user requirements. Constrained by space, we show the prompt case with one example in Figure 10.
@@ Appendix E Comparison Results of Model Responses
Appendix F Broader Impacts
The broader impacts of our work are as follows:
Enhancing Pediatric Medical Practice. Our model has made substantial contributions to pediatric medicine by integrating extensive medical data and emerging research. This integration facilitates more accurate and expedited diagnosis of complex pediatric conditions and aids in predicting treatment outcomes, enabling highly personalized and effective treatment strategies for young patients.
Medical Professionals and Public Health. Our PediatricsGPT provides crucial decision support for medical professionals, giving evidence-based recommendations and specialized medical insights. Additionally, it democratises access to expert medical suggestions and accurate medical knowledge, empowering parents and caregivers with accurate health information, which is especially crucial in underserved areas.
A Generalizable Training Framework for Broad Applications. The training pipeline of PediatricsGPT showcases exemplary generalizability, designed to be applicable across various medical and non-medical domains. This adaptability broadens the model’s applicability and pioneers the development of future AI solutions in healthcare and other fields.
Appendix G Limitations
The limitations of our work are as follows:
Security Risks of Model Manipulation. When deployed online, the proposed PediatricsGPT model, like other Large Language Models (LLMs), faces significant security risks, particularly from attacks aimed at manipulating its outputs. These attacks can be strategically designed to exploit the model’s response mechanisms, allowing attackers to induce the model to generate unsafe, biased, or otherwise inappropriate content.
Incomplete Language Coverage. Currently, the PediatricsGPT model does not support all languages. This linguistic barrier can prevent the model from reaching a global audience, particularly in diverse linguistic landscapes where localized medical information is crucial.
Appendix H Future Work

### v4

@@ Abstract
Developing intelligent pediatric consultation systems offers promising prospects for improving diagnostic efficiency, especially in China, where healthcare resources are scarce. Despite recent advances in Large Language Models (LLMs) for Chinese medicine, their performance is sub-optimal in pediatric applications due to inadequate instruction data and vulnerable training procedures. To address the above issues, this paper builds PedCorpus, a high-quality dataset of over 300,000 multi-task instructions from pediatric textbooks, guidelines, and knowledge graph resources to fulfil diverse diagnostic demands. Upon well-designed PedCorpus, we propose PediatricsGPT, the first Chinese pediatric LLM assistant built on a systematic and robust training pipeline. In the continuous pre-training phase, we introduce a hybrid instruction pre-training mechanism to mitigate the internal-injected knowledge inconsistency of LLMs for medical domain adaptation. Immediately, the full-parameter Supervised Fine-Tuning (SFT) is utilized to incorporate the general medical knowledge schema into the models. After that, we devise a direct following preference optimization to enhance the generation of pediatrician-like humanistic responses. In the parameter-efficient secondary SFT phase, a mixture of universal-specific experts strategy is presented to resolve the competency conflict between medical generalist and pediatric expertise mastery. Extensive results based on the metrics, GPT-4, and doctor evaluations on distinct downstream tasks show that PediatricsGPT consistently outperforms previous Chinese medical LLMs. The project and data will be released at https://github.com/ydk122024/PediatricsGPT.
@@ 2 Related Work
Chinese Large Language Model Evolution. The emergence of Large Language Models (LLMs) dominated by ChatGPT [33] and GPT-4 [5] has revolutionized the paradigm for novel human-machine interaction. Driven by learning-oriented technologies [11, 12, 13, 53, 54, 55, 48], pragmatic instruction [32, 47] and preference optimization [7, 34] strategies enable LLMs to address complex generation tasks with aligned human intentions. Despite improvements, large-scale resources for training general LLMs [28, 43, 44] are anchored in the English corpora, limiting their abilities to respond reliably in extensive Chinese application scenarios. Recently, researchers [18, 59] have attempted to enhance the comprehension and execution of Chinese instructions in open-source LLMs by augmenting Chinese vocabulary and data (e.g., Chinese LLaMA and Alpaca [18]). To facilitate Chinese-specific demands, several LLMs trained from scratch exhibit remarkable Chinese proficiency due to multilingual data resources, such as the Baichuan [2, 52], General Language Model (GLM) [20, 57], and Qwen [6] families. In this work, the Baichuan2-Base series is utilized as the foundation model for our PediatricsGPT, given its comprehensive potential among similar contenders.
@@ 3.1 PedCorpus: Multi-task Medical Instruction Dataset
Specialized Pediatric Data. Extracting pediatric data from textbooks, guidelines, and knowledge graphs ensures knowledge professionalism. Specifically, we automatically extract standard medical definitions and descriptions from physical textbooks covering 131 disease types in 11 broad categories. Over 500 corresponding disease guidelines are collected, including diagnostic protocols and treatment consensus. Additionally, extensive knowledge entities are sampled from ternary instances in the knowledge graphs. Based on these resources, we introduce a role-playing-driven instruction building rule via GPT-4 API that produces well-organized instructions to enable accurate and humanistic model responses. The detailed building procedure is shown in Appendix A.1.
@@ 3.2 Hybrid Instruction Pre-training in CPT
Continuous Pre-Training (CPT) is essential in developing domain-specific models [14, 49, 56] since it can break the scaling law [24] to a certain extent. For this purpose, we introduce the PedCorpus-CPT dataset to ensure a high-quality pre-training corpus. From Table 1, PedCorpus-CPT consists of three-part data components. (i) We integrate plain texts from vanilla pediatric textbooks, guidelines, and knowledge graphs. (ii) The filtered Chinese Wikipedia [3] is also considered to achieve the model’s trade-off for medical-general knowledge memory capacity. (iii) In practice, we observe that CPT leads to catastrophic forgetting of the models at follow-up due to different data distribution and format discrepancies compared to the original pre-training and SFT. Thus, we introduce a hybrid instruction pre-training mechanism to bridge these discrepancies. The core philosophy is to assemble instruction data from PedCorpus with Input-Output forms into Completion forms, which are then assimilated into plain texts to provide multi-task and complementary information. This mechanism effectively mitigates inconsistencies between the internal-injected medical knowledge of the foundation model while reinforcing medical domain adaptation. Moreover, we take PedCorpus as the seed instructions to improve multiple-department corpus density and breadth via knowledge-enhanced prompts. The prompt template is shown in Appendix B.
@@ 4.2 Model Zoo
We compare a series of LLMs for comprehensive evaluations. Concretely, Baichuan2-7B/13B (Chat) models [52] are trained on 2.6 trillion tokens as the baselines, which have excellent abilities in different domains. Meditron-7B [16] is a 7 billion parameters model adapted to the medical domain from Llama2-7B through continued pre-training on a comprehensively curated medical corpus. Llama3.1-8B [21] is a robust multilingual large language model through systematic training. For reproducible Chinese medical works, DISC-MedLLM (13B) [8] is fine-tuned through reconstructed medical dialogues and behavioural preference instructions. HuatuoGPT (13B) [58] performs SFT based on mixed instruction data and introduce human feedback in RLHF. HuatuoGPT-II (13B) [14] enhances the medical-specific domain adaptation of LLMs through one-stage unified training. Zhongjing (13B) [56] implements a complete pipeline based on Ziya-LLaMA-13B to enhance the model’s multi-turn medical conversation abilities. ChatGPT [33] and GPT-4 [5] have impressive performance in general medical fields as closed-source models developed by OpenAI.
@@ 4.4 Ablation Studies
Necessity of Supervised Fine-tuning. (i) We observe consistent performance gaps when removing the Full-parameter SFT (FSFT) and Parameter-efficient SFT (PSFT) phases, respectively. This makes sense since SFTs are necessary to activate the model’s healthcare instruction-following capabilities. (ii) Moreover, PSFT is more critical for three pediatric applications because it facilitates pediatric-related knowledge accumulation, while FSFT focuses on consolidating general medical semantic representations. (iii) Then, we replace the proposed Mixture of Universal-specific Experts (MUE) version with the vanilla single LoRA. The reduced performance on pediatric EviDiag and TreRecom benchmarks verifies that it is essential to introduce multiple LoRAs that act as specific experts on different tasks. A reasonable explanation is that the single-LoRA model suffers from the task competition between learning the knowledge question-answer and mastering the diagnostic recommendation abilities. (iv) Furthermore, we find that the universal LoRA expert significantly improves the results on the general medical benchmarks (i.e., CMD and webMedQA), proving that it mitigates the competency conflict between general medical and pediatric knowledge.
@@ 4.6 Visualization Analysis of Model Responses
To intuitively compare the output quality of medical LLMs, we show the responses of different models for each of the three types of medical inquiries in Figures 13&14&15 from Appendix E. From the results, Zhongjing offers insufficient information due to limited output content. Although HuatuoGPT-II gives well-organized responses, it lacks accuracy and informativeness. In comparison, our model can provide more specialized and detailed medical knowledge and diagnostic guidance in extended response contexts, confirming its application potential in diverse healthcare services.
5 Conclusion and Discussion
This paper presents PediatricsGPT, a Chinese medical LLM assistant with medical generalist and pediatric expertise capabilities. Based on the well-designed PedCorpus dataset, PediatricsGPT undergoes a systematic and robust procedure ranging from continuous pre-training and supervised fine-tuning to human preference optimization, leading to competence in different pediatric and general healthcare service scenarios. Extensive experimental results under multi-dimensional evaluation patterns demonstrate that our model outperforms currently available Chinese medical LLMs, providing a potential solution for promoting reliable and intelligent interactive diagnosis and treatment.
Broader Impacts. (i) Our model has made meaningful contributions to pediatric medicine by integrating extensive medical data and emerging research. This integration facilitates more accurate and expedited diagnosis of complex pediatric conditions and aids in predicting treatment outcomes, enabling highly personalized and effective treatment strategies for young patients. (ii) The proposed PediatricsGPT provides crucial decision support for medical professionals, giving evidence-based recommendations and specialized medical insights. Additionally, it democratizes access to expert medical suggestions and accurate medical knowledge, empowering parents and caregivers with accurate health information, which is especially crucial in underserved areas. (iii) The training pipeline of PediatricsGPT showcases exemplary generalizability, designed to be applicable across various medical and non-medical domains. This adaptability broadens the model’s applicability and pioneers the development of future AI solutions in healthcare and other fields.
Limitations. (i) When deployed online, the proposed PediatricsGPT model, like other Large Language Models (LLMs), faces significant security risks, particularly from attacks aimed at manipulating its outputs. These attacks can be strategically designed to exploit the model’s response mechanisms, allowing attackers to induce the model to generate unsafe, biased, or otherwise inappropriate content. (ii) Currently, our PediatricsGPT model does not support all languages. This linguistic barrier can prevent the model from reaching a global audience, particularly in diverse linguistic landscapes where localized medical information is crucial.
Ethical Issues. We fully recognize the critical importance of privacy and data protection. All data used has been meticulously de-identified, with all sensitive information removed, and this process has been verified by the partnering medical institutions. For the public databases, we strictly follow specific license agreements for use and adaptation. For the constructed corpus, we underwent an internal ethical review by the ethical review board of the partnering medical institutions with license and approval. We will release relevant resources to the extent that they are controlled and permitted.
We provide more discussions of the future work in Appendix F.
Acknowledgment. This work is supported in part by the National Key R&D Program of China under Grant 2021ZD0113502, in part by the Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0103.
@@ A.1 Role-playing-driven Instruction Building Rule
After integrating pediatric textbooks, guidelines, and knowledge graphs into a consolidated textual database, the content is segmented according to individual diseases. Subsequently, two instances of the GPT-4 model are deployed, designated as the “inquirer” and the “expert pediatrician” respectively. Disease-specific segments are then fed into the “inquirer” GPT-4, tasked with formulating a series of relevant and scholarly pediatric inquiries. Following this, the original disease segments and the formulated inquiries are fed into the “expert pediatrician” GPT-4 to generate precise responses for each inquiry, leveraging the segmented text as the contextual reference. We show the prompt templates for the “inquirer” and “expert pediatrician” in Figures 7 and 8, respectively.
@@ A.2 Vanilla Doctor-patient Conversation Regularization
We guide GPT-4 to regularize concise and noisy doctor responses in authentic doctor-patient consultations by the context learning strategy. Specifically, we manually craft 100 instruction examples with high-quality content to allow GPT-4 to learn doctor-like and patient-friendly behavioral styles. In each round of regularization, we randomly sample 10 out of 100 examples to perform 10-shot context prompts. Immediately after that, vanilla dialogues are fed to GPT-4 as seed instructions to optimize instructions according to user requirements. Constrained by the space, we show the prompt case with one example in Figure 9.
@@ Appendix E Comparison Results of Model Responses
Appendix F Future Work